OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to y from the first equation is substituted into the second equation. 8x - y = -9 4x - 3y = -22 (a)4(-8x - 9) - 3y = -22 (b)4x - 3(-8x - 9) = -22 (c)4(8x + 9) - 3y = -22 (d)4x - 3(8x + 9) = -22

OpenStudy (anonymous):

IN A HUGE HURRY. ANSWER PLEASEEEE.

OpenStudy (unklerhaukus):

To perform a substitution, you have to isolate of of the variables, the y in the first equation is the easiest to isolate what is y = ?

OpenStudy (unklerhaukus):

to isolate y from the equation 8x - y = -9 add y to both sides and also add 9 to both sides what do you get?

OpenStudy (unklerhaukus):

. . .

OpenStudy (unklerhaukus):

@ElijahM

OpenStudy (anonymous):

Im lost....

OpenStudy (anonymous):

y=9?

OpenStudy (unklerhaukus):

8x - y = -9 adding y to both sides 8x - y + y = -9 + y the left hand side simplifies 8x = -9 + y now add 9 to both sides 8x +9 = -9 + y + 9 now the right hand side simplifies . . .

OpenStudy (anonymous):

soooo 8x+9=y would be it?

OpenStudy (unklerhaukus):

good work you have isolated y; (y = 8x+9) now look at the second equation 4x - 3y = -22 substitute (y = 8x+9) into this equation what do you get ?

OpenStudy (anonymous):

4x-3(8x+9) being answer D?! (:

OpenStudy (unklerhaukus):

dont forget the [= - 22] part but Great work ! you've done it \(\color{red}\checkmark\)

OpenStudy (anonymous):

Thank you Thank you! (: I understand now lol

OpenStudy (unklerhaukus):

\[\ddot\smile\]